Question 15

  1. N2O (laughing gas) is used as an anesthetic. A dentist has a 2.00 L flask of N2O at a pressure of 1.00 atm and a 3.00 L flask of nitrogen gas at 2.00 atm that are separated by a valve. What is the pressure in the flasks when the valve is opened, and the gases mix together? Both flasks are at the same temperature, and the temperature does not change when the gases mix.
Answer #1

We will use Boyle’s law which is:

Pi*Vi = Pf*Vf

P1*V1 + P2*V2 = Pf*Vf

P1*V1 + P2*V2 = Pf*(V1+V2)

1.00*2.00 + 2.00*3.00 = Pf*(2.00+3.00)

8.00 = Pf*5.00

Pf = 1.60 atm

Answer: 1.60 atm